OpenStudy (anonymous):

2x + 5 = 10 What is the first operation used in solving this equation? A) addition B) division C) multiplication Eliminate D) subtraction

OpenStudy (landon34):

2x + 5 = 10 2x = 10 - 5 2x = 5 (divide both sides by 2 to get x) 2x/2 = 5/2 x = 2.5

OpenStudy (landon34):

thats the entire problem solved, you are welcome.